Problem 1. Consider the harmonically forced undamped oscillator described by the following ODE: mx + kx = Fo cos wt, k > 0,Problem 1.Consider the harmonically forced undamped oscillator described by the following ODE:mx′′+kx=F0cosωt, k >0, m >0, ω >0, F0∈R.
Answer #1

Q1. Sok Given data: mx + &x = to cosut, - ko, no, w OFER. The given ODE can be written as x + ku= to cos(uf) - . (2) Given th(6) IVP: So x(0)= 0 and (0) = 0 G coso + C sinot to coso o at tao. -> 4+ [ cos 0 1 Sinozo 6] → 9 -ws) Again n (t) = - Wo G sComparing en ④ with a(t) = (sin (wit) sin ust) we get wy = w two and. Wq = www.

The problem is solved and provided above.

Here, I've used a standard result to find the particular solution, i.e. \phi_p=\frac{1}{f(D)} f(t) where F(D)= D^2+\omega_0, D= \frac{d}{dt} and f(t)= \frac{F_0}{m}cos(\omega t) , otherwise it'll be quite lengthy.

Method of variation of parameters can also be used to find it, and result will be same.

Thank you.